Wednesday, 22 December 2010

Harsh Winter

The weather did go milder for a week, I took the monitoring boards out, but then went cold again (minus double figures, freezing during the days). The boards went back in. The four large colonies have liquid on their monitoring boards, it is around the edge ( not where the colony is). This may be because of how the condensation collects or any drips onto the bees are dispersed. Photo of one of the boards.

Monday, 6 December 2010

Early winter

Placed monitoring board under the small colony last week and all other colonies today.It's been so cold, minus 12 this morning, minus 10 tonight. Weather like this for over a week!

Sunday, 14 November 2010

Winter prep

I placed an empty super under the brood box last week, and this week I have removed the monitoring floor.
I hope this will make their nest cooler which will result in no brood and less stores consumed. The colony that was created late is not very big and so I will replace the monitoring floor if it gets very cold!

Sunday, 10 October 2010

even more

more images

Great weather

This afternoon it is 20+ degrees, lots of activity and bringing in pollen at the hives, and lots of butterflies on the Michaelmas Daisy's

Tuesday, 28 September 2010

Still flying

Still bringing pollen in today

Saturday, 18 September 2010


Compared with last year the Varroa drop is quite low, at most 20 per day. So I have decided not to give a second treatment to any of the colonies.
THe first colony was very low on bees, which does not look good for winter. I will need to give it some insulation in deep winter. I wonder if the Apiguard has made some abscond?
I checked to find only a tiny area of brood, I wonder if the Apiguard has stopped her laying.
THe first and forth had not taken down any more feed so I removed the feeders and gave the syrup to the others.

Sunday, 12 September 2010

Apiguard 2

The drop over this week has not been very high so decided the crown board may have been limiting the spread, sealing against the foil sachets. I turned the sachets over which will give colony high exposure. Today there was a high drp from all hives.
On the colonies I did look into there seems a lot of stores close to the brood, don't think I will feed for a while, so far they have had 30Kg between them

Friday, 3 September 2010


I have put a Apiguard sachet on every colony this afternoon, if some of the colonies still have low drop I may not use a second sachet.

Tuesday, 31 August 2010

Wednesday, 25 August 2010


Extracted today, loaned spinner again. Got 12kg this year only 5kg last. Very messy, need a better way to drain cappings (they would not spin in mesh bag as they touched side), the honey would not pass through the fine strainer.

Sunday, 22 August 2010

Start weighing

The roofs and ekes are different weights.
Estimating each hive, the following are the weights to deduct to work out stores in colony. THese include roofs and ekes if they have them.
1 - 17.36kg
2 - 16.8kg
3 - 17.3kg
4 - 19.54kg
5 - 22.68kg
The estimatd stores are on the colony spreadsheets.

Saturday, 17 July 2010


The second colony had produced QC just over a week ago, so left queen in new box with 3 frames of comb and the rest starter strips, plus supers (artificial swarm). All the brood and qc's moved to differnt position. THe swarm have built new comb on 5 frames very quick, looking good.

Thursday, 24 June 2010

Another Nuc

Made another split tonight, 3 frames with brood on, two QC, two frames of stores and space. Left my original colony with 8 frames and QC.
Shown is my first analysis of my original colony, 30 wings.

Monday, 21 June 2010


There were open QC on 2 frames on Saturday so need to make up the nucs no later than next Saturday, don't know wether to do 2 or 3 yet.. Every time I touch the queenless colony the bees race out, very lively! Collected around 30 bees from it tonight to anylise their wings, they were very noisy as they entered the freezer.

Wednesday, 16 June 2010


Opened my new nuc tonight as I wont be able to tomorrow. Shook out the last failed nuc and all the bees tryed to enter the new nuc, and they fought! Need to find a better way to empty a failed nuc, hope my original queen is OK.

Tuesday, 15 June 2010

Try Again

Took my original queen (white) from my main hive with 2 frames with brood, 2 frames with stores and pollen, 1 empty comb. Will seal for 2 days and then place where my last failed nuc is (position 4), where it can pick up returning bees. Bees were robbing empty hives tonight, should have been more careful! The original hive is strong and will now make a emergency queen, if there are spare cells I may try another nuc from the other hive. There should be a queen cell capped in 5/7 days and an emerging queen in 12/14 days.

Sunday, 13 June 2010

bad news

Checked the nucs today, they all seem to have failed. One QC has died and in the other nuc brood is hatching small, black and wet looking, I think the flying bees have returned to their original colony and the brood has chilled. Back to the drawing board.

Sunday, 6 June 2010

New Splits

Checked the 3 QC on one frame today, one had emerged. So created 2 nucs with a frame of mainly sealed brood, 2 frames with pollen and stores and a QC. The emerged queen may have been on the remaining frames, so left in the brood box and placed in the following pattern on the QC raising colony.
Original colony in brood box,
Queen excluder
2 supers.
Queen excluder.
crown board with side opening and bee proof mesh over holes
Queen cell raising brood box.
This would allow the new queen to mate and start laying in the top brood box and making a 2 queen colony.
I'm sure the nuc created two weeks ago is queenless, so if I get the chance I will place a queen cell with it, also because I don't know if the emerged queen is still with the top box I would give another QC.

Sunday, 30 May 2010

Having a drink

Queen Check

Queen cells still uncapped today and lava don't look too big, I think it's 8 days after capping!

Nuc with John's QC, both cells are empty, just had a quick peek and left.

Friday, 28 May 2010

Queen rearing`

Made original hive queenless 4 days ago, gave frame with young lave a day later. Today reunited colony, 4/5 QC which appear small but are uncapped. need to check in a few days

Sunday, 23 May 2010

New nuc

Placed johns QC in on Friday night, was going to leave them sealed for 5 days somewhere cool but yesterday the were buzzing so much I moved them to bottom of garden and opened. Weather has been very hot, late twenties

Thursday, 20 May 2010

First Q cells

John phoned today, 7 QC, and he offered me one. Made a small nuc tonight from the middle one,
one frame with sealed brood with a frame with stores on either side. The warm way, shook a few more bees in, sealed them up and placed somewhere cool ( the weather at the moment is very hot, 20+) put a feeder on.

Sunday, 16 May 2010

TReating varroa

Turned over the Apigaurd to expose more thymol to the brood yesterday, big drop today. MOved last brood above excluder yesterday ready for Queen rearing.

Tuesday, 11 May 2010

Cold weather

Even though weather is poor and day temp has not reached 10 degrees today the Apiguard is causing big mite fall. Need to get ready for queen production and nucs.

Tuesday, 4 May 2010

Another check

Right colony is a lot bigger than left, both a lot slower build up compared to last year.
There were a few varroa on monitoring boards so will now check over a week. New comb seems very yellow!
I am going to turn colonies so all will get morning sun and I can still access monitoring boards.

Saturday, 24 April 2010

Quick check

Weather still lovely, frost first thing, 15+ by afternoon. Colonies seem alright, still introducing starter frames on edge of brood. First frame of drone brood in right colony.

Saturday, 10 April 2010

Quick Check

Nice warm day, lots of activity. Both colonies have built around a third of a frame of new comb, removedentrance blocks. The photo is not as good as I would like but:_

Sunday, 4 April 2010

Still cool

Quick check, bees built negligible comb on the starter strips and only taken a little syrup down. Used Cheddleton Weather station to work out average last 7 days = 5.45 degrees. I obviously started too early, must check when things start to move.

Saturday, 27 March 2010

Comb replacement

Put frames with starter strips either side of frames with brood, mixed 4kg with 4 lts, started feeding with up and over feeders. The monitoring boards will help to keep the colony warm, maybe if the colonies appear to have plenty of stores I should replace boards earlier than I did. This would help kick start brood rearing.

Sunday, 21 March 2010

Colony active

This last week the weather has warmed up 10 degrees plus, frogspawn has arrived. Both colonies were very active today, there were a cup full of dead bees on the floor of each hive and both have plenty of stores. I have removed the empty super from under the brood, put monitoring boards in and entrance blocks in. They were both bringing in plenty of pollen, two colours?

Sunday, 14 March 2010

Exact Weights

Measuring the colonies by weighing each side is very inaccurate! (1 or 2 KG out).
So I measured all the seperate items from the deceased colony in KG.
Crown .96
Super box 3.68
Brood box 3.26
Mesh floor & block 1.4
Frame & comb, no stores, approx .25
Homemade roof 5
Bought roof 4.5
Floor, Super, Brood, frames, crown 12
There was 1.6 KG of stores still in the hive.
My estimation of 2.36 for a super was wrong, and so this winters calculations are out. If each item always weights the same my currant winter setup (floor, super, brood, frames, crown) =12Kg. I guess 10,000 bees weights around 1Kg.
Next year for a full size colony, if I take 13Kg from it's weight I will have an estimation of their stores.

Saturday, 6 March 2010

First check

THe small left colony has died. There was around 8lb of stores scaterred around the frames so I presume it was cold. The cluster was only small and only used the a single space between two frames. There was no brood and I did not see any pollen. This was my smallest colony, it had made an emergancy queen during August.
Gave the other two some fondant.